OpenStudy (anonymous):

How would I get the graph for this and what would it look like? g(x)= e^((−3x+2sinx)/6) ------------------ (3+2cosx)

OpenStudy (anonymous):

What, again. Simplest is to feed it to Wolfram. Else note that it always positive, calculate maxima/minima, end behaviours etc

OpenStudy (anonymous):

I have done that but I need to understand how i get to the answer through calculation. Could you guide me through it.

OpenStudy (anonymous):

I am a bit lost on this problem

OpenStudy (anonymous):

You can see that the top and bottom are positive so you need not concern yourself with anything below the x-axis.

OpenStudy (anonymous):

I forget what they were but you get the derivatives and possible min/max (then test either side for increasing/decreasing or else use second derivative to check if min or max.

OpenStudy (anonymous):

Now, what happens if x= 0?

OpenStudy (anonymous):

Now consider what happens for increasing negative and positive vales of x.

OpenStudy (anonymous):

Sketch it.

OpenStudy (anonymous):

Do a couple more values if necessary.

OpenStudy (anonymous):

Oh and I forgot, there is a value of x that makes the denominator zero, need to look at that as well.

OpenStudy (anonymous):

I get to this point but can't make sense of what they mean in terms of the graph. (sinx -5/2) and (sinx-1/2). These are meant to be the stationary points. But I can visualise it.

OpenStudy (anonymous):

I don't understand what you mean? How can sinx -1/2 be a stationary point?

OpenStudy (anonymous):

I mean stationary points are when x is 1/2 and x is 5/2.

OpenStudy (anonymous):

OK, I assume that those are correct, now we need to know whether they are max or min (you can test either side for positive/negative or use second derivative)

OpenStudy (anonymous):

You then need to calculate y at the min/max to be able to plot those points.

OpenStudy (anonymous):

Do they seem right to you? I will now test those point.

OpenStudy (anonymous):

Wait a minute, I will get a graph from Wolfram to check.....

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=plot+ [g%28x%29%3D++e^%28%28%E2%88%923x%2B2sinx%29%2F6%29+%2F%283%2B2cosx%29%2C{x%2C-10%2C5}%2C{y%2C-1%2C10} Um, can you check if I have the correct equation, I am seeing max/min somewhere else

OpenStudy (anonymous):

Paste plot [g(x)= e^((−3x+2sinx)/6) /(3+2cosx),{x,-10,5},{y,-1,10} into Wolfram box

OpenStudy (anonymous):

plot [g(x)= e^((−3x+2sinx)/6) /(3+2cosx),{x,-10,5},{y,-1,10}]

OpenStudy (anonymous):

Here it is

OpenStudy (anonymous):

OK, I found it now..

OpenStudy (anonymous):

Ok, the equation is correct at Wolfram. So where did you get 1/2 and 5/2 from?

OpenStudy (anonymous):

I got it from working out 4sinx^2 +12sinx -5 =0 I used factoring and quadratic formula to get the stationary points

OpenStudy (anonymous):

4sinx^2 +12sinx + 5 =0 not -5

OpenStudy (anonymous):

And I don't get 1/2 and 5/2

OpenStudy (anonymous):

is it not -4sinx^2 -12x +5 or 4sinx^2 +12sinx -5. Are they not the same?

OpenStudy (anonymous):

Above, you said "4sinx^2 +12sinx -5 =0" it is +5 not -5 It also says 4sinx^2 +12sinx -5 =0 in the doc you sent me.

OpenStudy (anonymous):

It also says 4sinx^2 +12sinx +5 =0 in the doc you sent me.

OpenStudy (anonymous):

Anyway, this is a periodic function

OpenStudy (anonymous):

there is a minus sign sign just after g'(x). What happens to that?

OpenStudy (anonymous):

Maybe that is why I am getting a bit lost. Because the exercise up to now have not given me any exercise in this type of problem.

OpenStudy (anonymous):

If you want to you can absorb it into the eqaution to get -4sin^2 x +12 sin x -5

OpenStudy (anonymous):

Anyway, here is Wolfram plot[4sinx^2 +12sinx +5 =0 ,{x,-10,10},{y,-10,10}]

OpenStudy (anonymous):

can I just clarify something

OpenStudy (anonymous):

Now you have me doing it! plot[4sinx^2 -12sinx +5 =0, {x,-10,10},{y,-10,10}] is correct.

OpenStudy (anonymous):

I'm sorry, I really do not want to do any more of these computations.

OpenStudy (anonymous):

\[- ((4sinx-12sinx +5)\exp((-3x+2sinx)/6)/(6(3+2cosx)^2)\]

OpenStudy (anonymous):

−((4sinx−12sinx+5)

OpenStudy (anonymous):

It is sin^2 x not sin x

OpenStudy (anonymous):

Is this what you have in the file I sent you for the denominator?

OpenStudy (anonymous):

(4 sin^2 x -12 sin x +5) and there is a minus sign in front of the whole derivative which you can absorb anywhere you want to (or not).

OpenStudy (anonymous):

Anyway, we are spending too much time and getting nowhere, so I have to go now, i am sorry.

OpenStudy (anonymous):

Ok. Thanks for your help

OpenStudy (anonymous):

OK, I hope you can solve it (it really is not too difficult once you concentrate on it)